2. A vector A has a magnitude of 50.0 m and points in a direction 28.0° above the negative x-axis. A second vector, B. has a magnitude of 80.0 m and points in a direction 60.0 above the negative x-axis. Using the component method, find the magnitude of the vector R- Ä + B. x Blank m 3. A vector A has a magnitude of 50.0 m and points in a direction 28.0° above the negative x-axis. A second vector, B, has a magnitude of 80.0 m and points in a direction 60.0° above the negative x-axis. Using the component method, find the magnitude of the vector D = A-B XBlank m
Answer #1

angle between A and B

x = 60- 28 = 32

R^2 = A^2 + B^2 + 2 A B cos x

R^2 = 50^2 + 80^2 + 2* 50*80* cos 32

R = 125.24 m

======

A = 50 ( - cos 28 i + sin 28 j) = - 44.147 i + 23.47 j

B = 80 ( - cos 60 i + sin 60 j) = - 40 i + 69.282 j

D = A - B = ( - 44.147 + 40) i + ( 23.47 - 69.282)j

D = - 4.147 i - 45.812 j

magntidue

D^2 = 4.147^2 + 45.812^2

D = 46 m

=======

do comment in case any doubt, will reply for sure. Goodluck